40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 8 11:41:16 2018, gautam, Update, General, IFO maintenance ITMX_stuck.png
    Reply  Wed May 9 17:30:04 2018, gautam, Update, General, Input beam misaligned InputBeamGone.pngTTpointing.png
       Reply  Wed May 9 19:51:07 2018, gautam, Update, General, Input beam misaligned DACweirdness.pngDACerror.png
          Reply  Sun May 13 15:15:18 2018, gautam, Update, General, CDS crash vertexFEs_crashed.png
             Reply  Sun May 13 17:31:51 2018, gautam, Update, General, CDS crash CDS_overview_20180513.pngAS_1210293643.jpeg
                Reply  Sun May 13 20:48:38 2018, johannes, Update, General, CDS crash 
                Reply  Thu May 17 11:56:37 2018, gautam, Update, General, EPICS process died on c1ioo 
                   Reply  Thu May 24 10:16:29 2018, gautam, Update, General, All models on c1lsc frontend crashed c1lsc_crashed.png
    Reply  Thu May 10 08:45:16 2018, Steve, Update, General, 4.5M eq. Cabazon, CA Cabazon4.5m79m.png4.5Meq.png
       Reply  Thu May 10 11:38:19 2018, gautam, Update, General, ITMY UL ITMY_UL.pdf
       Reply  Wed Aug 22 08:44:09 2018, Steve, Update, General, earth quake yesterday_EQs.png
          Reply  Fri Aug 24 08:04:37 2018, Steve, Update, General, small earth quake small_EQ.png
             Reply  Mon Aug 27 09:14:45 2018, Steve, Update, PEM, small earth quakes  small_EQs_vs_SUSs.png
                Reply  Tue Nov 27 10:50:20 2018, Steve, Update, PEM, earth quake Mexico 5.5M.Mexico.png
                   Reply  Thu Nov 29 08:13:33 2018, Steve, Update, PEM, EQ 3.9m So CA 3.9mSoCA.pngVac_as_today.pngas_is.png
          Reply  Wed Aug 29 09:20:27 2018, Steve, Update, SUS, local 4.4M earth quake 4.4_La_Verne.png3.4_&_4.4M_EQ.png
             Reply  Wed Aug 29 11:46:27 2018, Jon, Update, SUS, local 4.4M earth quake 
             Reply  Thu Sep 20 08:17:14 2018, Steve, Update, SUS, local 3.4M earth quake local_3.4M.png
Message ID: 13827     Entry time: Wed May 9 17:30:04 2018     In reply to: 13826     Reply to this: 13828
Author: gautam 
Type: Update 
Category: General 
Subject: Input beam misaligned 

There is no beam going into the IFO at the moment. There was definitely a spot on the AS camera after I restored the suspensions yesterday, as you can see from the ASDC level in Attachment #1. But at around 2pm Pacific yesterday, the ASDC level has gone to 0. I suspect the TTs. There is no beam on the REFL camera either when PRM is aligned, and PRM's DC alignment is good as judged by Oplev.

Normally, I am able to recover the beam by scanning the TTs around with some low frequency sine waves, but not today. We don't have any readback (Oplev/OSEM) of the TT alignment, and the DC bias values havent jumped abnormally around the time this happened, judging by the OUT16 monitor points (see Attachment #2). The IMC was also locked at the time when this abrupt drop in ASDC level happened. Unfortunately, we don't have a camera on the Faraday so I don't know where the misalignment is happening, but the beam is certainly not making it to the BS. All the SOS optics (e.g. BS, ITMX and ITMY) are well aligned as judged by Oplev.

Being debugged now...

Attachment 1: InputBeamGone.png  43 kB  Uploaded Wed May 9 18:36:24 2018  | Hide | Hide all
InputBeamGone.png
Attachment 2: TTpointing.png  36 kB  Uploaded Wed May 9 18:41:15 2018  | Hide | Hide all
TTpointing.png
ELOG V3.1.3-